OpenStudy (anonymous):

Which is 243 written as a power of a number? A. B. C. D.

OpenStudy (anonymous):

@satellite73

OpenStudy (the_fizicx99):

._. you might wanna look at your options >.>

OpenStudy (anonymous):

I can't see what your choices A - D are, but a cheeky answer is 243 = 243^1.

OpenStudy (anonymous):

Here's a hint, though. 243 is a multiple of 3.

OpenStudy (anonymous):

oh crap sorry lol

OpenStudy (the_fizicx99):

3^5 = 243 >_<

OpenStudy (anonymous):

here

OpenStudy (the_fizicx99):

I'm good :O

OpenStudy (anonymous):

Yes you are XD ~ Thankies ~

OpenStudy (mathmale):

First, why not do a relatively simple experiment? Realizing that 243 is an odd number, try odd bases with positive integer exponents to determine whether or not you can arrive at 243. \[5^{1}=5; 5^{2}=25; 5^{3}=125; and so on.\] \[3^0=1; 3^1=3; 3^2=9; 3^3=27; and. so .on. \] What happens as you continue to increase the exponents of 3?

OpenStudy (the_fizicx99):

He covered it all >_< ^

OpenStudy (mathmale):

Later, you might recognize that 243=(3^2)(3^3) = 3^5.

OpenStudy (anonymous):

Learning the basic tests for divisibility helps too. I knew that 243 was a multiple of 3 because the digits of the number add to a multiple of 3, i.e. 2+4+3 = 9 = 3x3.

OpenStudy (mathmale):

What a neat suggestion, Cliff!

OpenStudy (mathmale):

Where'd you pick up that trick?

OpenStudy (the_fizicx99):

I usually look at the back of the number and see what its divisible by

OpenStudy (anonymous):

4th grade, I think . .

OpenStudy (anonymous):

I discovered a few others on my own. I now know how to test for divisibility of all numbers up to and including 12. I understand the general method for testing prime numbers greater than 12, but it's pretty complicated.

OpenStudy (anonymous):

Here's a cool one for divisibility by 11. The sum of every-other digit minus the sum of the other digits is a multiple of 11. e.g. 12705 - take 1+7+5 and subtract 2+0 and 13-2 is 11, so 12705 is divisible by 11.

OpenStudy (the_fizicx99):

What about 5,896,547 o.0

OpenStudy (anonymous):

It's not divisible by 3 or 11 (or any even number, of course). Not divisible by 5.., Doesn't appear to be divisible by 7 either, so I'd have to test some two-digit prime numbers next

OpenStudy (anonymous):

5896547 is divisible by 43.

OpenStudy (mathmale):

Cliff, I'm really impressed. I've made and am saving copies of this dialogue for later ref. Thanks.

OpenStudy (anonymous):

If you get a chance to study some number theory, especially modular arithmetic, it'll make more sense.

OpenStudy (anonymous):

Oh, and 5896547 is also divisible by 241 and 569.

OpenStudy (mathmale):

I really do need to express my reactions to those last few statements: WOW!

OpenStudy (the_fizicx99):

Dayummm >_<

OpenStudy (anonymous):

In the spirit of full disclosure, I found 241 and 569 by brute force, not any number theory trick.

OpenStudy (the_fizicx99):

Brute force is a program I used to decode games and modify them >.> didn't know that

OpenStudy (anonymous):

The term, "brute force" just refers to using the fast processing speed of a computer to grind out calculations. I wrote a quick program to iterate factor-finding calculations until it hit on a couple.

OpenStudy (the_fizicx99):

ah